OpenStudy (anonymous):

Clear the parentheses and combine like terms. -(2m - 7n) - 3(6m - n) Show steps, please! :)

OpenStudy (chaise):

-(2m - 7n) - 3(6m - n) Expand your negative from the front of the parantheses using the distributive law. -1(2m - 7n) - 3(6m - n) = -2m+7n+(-18m) + 3n = -2m+7n-18m+3n = -2m-18m+7n+3n = -20m+10n = 10n-20m Sorry if this has confused you, let me know an I will try and clear up any confusion.

OpenStudy (anonymous):

No confusion, thank you! :)